# Learning Contrastive Feature Representations for Facial Action Unit Detection

### v1

@@ IV-A3 Evaluation Metrics
V Evaluation Metrics
TABLE I: Quantitative comparison results on BP4D dataset, where the evaluation metric is F1-score.
@@ V Evaluation Metrics
KDRL [63] 2022 CVPR 53.3 47.4 56.2 79.4 80.7 85.1 89.0 67.4 55.9 61.9 48.5 49.0 64.5
AUNet [64] 2023 Proc.IEEE 58.0 48.2 62.4 76.4 77.5 83.4 88.5 63.3 52.0 65.5 52.1 52.3 65.0
AUNCE(Ours) - 54.3 49.5 60.8 78.3 78.1 84.4 89.1 66.6 56.3 64.8 49.6 56.5 65.7
TABLE II: Quantitative comparison results on DISFA dataset, where the evaluation metric is F1-score.
@@
MAL [58] 2023 TAC 43.8 39.3 68.9 47.4 48.6 72.7 90.6 52.6 58.0
SEV-Net [61] 2021 CVPR 55.3 53.1 61.5 53.6 38.2 71.6 95.7 41.5 58.8
HMP-PS [59] 2021 CVPR 38.0 45.9 65.2 50.9 50.8 76.0 93.3 67.6 61.0
MEGraph [51] 2022 IJCAI 52.5 45.7 76.1 51.8 46.5 76.1 92.9 57.6 62.4
@@
SupHCL [62] 2022 ACM MM 52.5 58.8 70.0 53.5 51.4 73.1 95.6 58.0 64.1
AAR [60] 2023 TIP 62.4 53.6 71.5 39.0 48.8 76.1 91.3 70.6 64.2
KDRL [63] 2022 CVPR 60.4 59.2 67.5 52.7 51.5 76.1 91.3 57.7 64.5
MMA-Net [41] 2023 PRL 63.8 54.8 73.6 39.2 61.5 73.1 92.3 70.5 66.0
AUNet [64] 2023 Proc.IEEE 60.3 59.1 69.8 48.4 53.0 79.7 93.5 64.7 66.1
AUNCE(Ours) - 61.8 58.9 74.9 49.7 56.2 73.5 92.1 64.2 66.4
V-A Comparison with State-of-the-Art Methods
While our discriminative contrastive learning primarily operates as a self-supervised framework, our experimental results on BP4D and DISFA still compare with some supervised frameworks, because we use AU labels assisting the process of selecting positive sample pairs. The focus of our comparison lies in evaluating state-of-the-art methods from the past three years, outlined as follows:
  • HMP-PS (CVPR 2021) [59] presents a pioneering approach to message passing neural networks incorporating performance-driven structures.
  • SEV-Net (CVPR 2021) [61] proposes a strategy for AU detection by leveraging semantic embedding and visual features.
@@ V-A Comparison with State-of-the-Art Methods
  • SupHCL (ACM MM 2022) [62] introduces a method for AU detection that utilizes supervised hierarchical contrastive learning to ensure knowledge consistency across different facial images and AUs.
  • KDRL (CVPR 2022) [63] employs AU labeling rules from the Facial Action Coding System for self-supervised representation learning in AU detection.
  • MEGraph (IJCAI 2022) [51] focuses on deep learning of a unique graph to explicitly describe relationships between AUs in the target facial display.
  • MMA-Net (PRL 2023) [41] proposes a novel AU detection framework based on multiview mixed attention.
  • AUNet (Proc. IEEE 2023) [64] presents AU detection coupled with facial landmark detection, a simple yet robust baseline for landmark-based AU detection.
  • MAL (TAC 2023) [58] employs a meta learning approach to automatically select highly related facial expression samples using adaptative weights for training facial expression samples in AU detection.
  • AAR (TIP 2023) [60] introduces an adaptive attention regression network to regress the global attention map of each AU, guided by attention predefinition and AU detection.
  • CLP (TIP 2023) [23] aims to alleviate data scarcity issues by learning AU representations from a large pool of unlabeled facial videos in a contrastive learning paradigm.
The quantitative comparative results for the BP4D dataset are outlined in Table I. AUNCE exhibits superior performance in overall AU detection when contrasted with alternative approaches.
In comparison to MMA-Net, HMP-PS, AAR, SEV-Net, and KDRL, our approach exhibits a performance advantage of approximately 1.2%-2.3% in F1-score. This underscores the efficacy of our supervised discriminative contrastive learning paradigm in effectively capturing pertinent information for the training process. When contrasted with other contrastive learning methods such as CLP and SupHCL, our method demonstrates an average improvement of 3.3% and 1.3%, respectively. This improvement is attributed to our positive sample sampling strategy, which alleviates issues associated with AU labels tainted by noise and errors, enabling the encoder to extract differences among static image features more comprehensively.
@@
Compared to the second-best approach, MEGraph, our method demonstrates superior performance by elevating the F1-score from 65.5% to 65.7%. It is noteworthy that the encoder used in our method is only the same as the network MEGraph used for the first stage of training, with reduced parameters by nearly half. This highlights the applicability of AUNCE to more lightweight network structures and emphasizes that learning only the difference information among features is more potent than learning the pixel-level information of the entire image.
Comparable outcomes are evident in the DISFA dataset, as outlined in Table II, where AUNCE attains the highest overall performance in AU detection.
V-B Ablation Study
TABLE III: Quantitative results of ablation study on BP4D dataset.
Model BL PS IR F1-Score(%) Accuracy(%)
A $\checkmark$ 60.1 79.1
B $\checkmark$ $\checkmark$ 63.2(+3.1) 79.7(+0.6)
C $\checkmark$ $\checkmark$ 63.4(+3.3) 79.8(+0.7)
D $\checkmark$ $\checkmark$ $\checkmark$ 65.7(+5.6) 80.2(+1.1)
V-B1 Ablation Study
The quantitative results of our method with various component combinations are summarized in Table III, encompassing the baseline (BL) rooted in the new discriminative contrastive learning paradigm, utilizing images with the highest similarity to original images as positive sample pairs, the positive sample sampling strategy (PS), and the importance re-weighting strategy to minority AUs (IR). The merits of our proposed discriminative contrastive learning paradigm will be elaborated in Section V-B2.
Contribution of the positive sample sampling strategy: As outlined in Table III, the positive sample sampling strategy contributes to a 3.1% increase in F1-score and a 0.6% improvement in Accuracy compared to the baseline. This validates that selecting different types of positive sample pairs with appropriate probabilities can mitigate issues arising from AU labels tainted by noise and errors. The strategy achieves this by incorporating three types of positive sample pairs: the self-supervised signal pair, the highest similarity pair, and the mixed positive features pair. These pairs provide the models with diverse information from various views, thereby enhancing the generalization ability of the contrastive learning framework.
Effectiveness of the importance re-weighting strategy: The proposed importance re-weighting strategy to minority AUs results in a 3.3% increase in F1-score compared to model A. This illustrates that the strategy can change the backpropagation gradient of positive and negative samples, facilitating faster label category gradient updates with a smaller number of samples. Additionally, model D achieves an average increase of 2.5% in F1-score compared to model B, further supporting the efficacy of the importance re-weighting strategy.
Based on ablation studies, the introduction of each component proves beneficial for performance improvement, with IR emerging as the most effective component.
@@ V-B1 Ablation Study
TABLE IV: Comparison of the weighted cross-entropy loss and AUNCE on BP4D and DISFA.
Dataset Loss MEGraph-stage1
F1-Score Accuracy
BP4D WCE 63.6 80.0
AUNCE 65.7(+2.1) 80.2(+0.2)
DISFA WCE 62.2 93.7
AUNCE 66.4(+4.2) 94.3(+0.6)
V-B2 Compared with weighted cross-entropy loss

### v3

@@ IV-A3 Evaluation Metrics
TABLE II: Quantitative comparison results on BP4D dataset, where the evaluation metric is F1-score(%).
@@
ISTR [80] 2021 ACM MM 54.0 46.0 55.7 79.4 78.8 84.5 87.0 67.0 55.6 63.1 50.7 55.3 64.8
SimCLR [22] 2020 ICML 38.0 36.4 37.2 66.6 64.7 76.2 76.2 51.1 29.8 56.1 27.5 37.7 49.8
MoCo [32] 2020 CVPR 30.8 41.3 42.1 70.2 70.4 78.7 82.5 53.3 25.2 59.1 31.5 34.3 51.6
BTFM [81] 2023 CVPR 57.4 52.6 64.6 79.3 81.5 82.7 85.6 67.9 47.3 58.0 47.0 44.9 64.1
@@
KSRL [46] 2022 CVPR 53.3 47.4 56.2 79.4 80.7 85.1 89.0 67.4 55.9 61.9 48.5 49.0 64.5
CLEF [82] 2023 ICCV 55.8 46.8 63.3 79.5 77.6 83.6 87.8 67.3 55.2 63.5 53.0 57.8 65.9
AUNCE(Ours) - 53.6 49.8 61.6 78.4 78.8 84.7 89.6 67.4 55.1 65.4 50.9 58.0 66.1
(±$\pm 0.15$) (±$\pm 0.12$) (±$\pm 0.17$) (±$\pm 0.21$) (±$\pm 0.19$) (±$\pm 0.22$) (±$\pm 0.22$) (±$\pm 0.17$) (±$\pm 0.14$) (±$\pm 0.15$) (±$\pm 0.14$) (±$\pm 0.15$) (±$\pm 0.17$)
TABLE III: Quantitative comparison results on DISFA dataset, where the evaluation metric is F1-score(%).
@@
EAC-Net [7] 2018 TPAMI 41.5 26.4 66.4 50.7 80.5 89.3 88.9 15.6 48.5
ROI [4] 2017 CVPR 41.5 26.4 66.4 50.7 80.5 89.3 88.9 15.6 48.5
ARL [15] 2019 TAC 43.9 42.1 63.6 41.8 40.0 76.2 95.2 66.8 58.7
MAL [75] 2023 TAC 43.8 39.3 68.9 47.4 48.6 72.7 90.6 52.6 58.0
[…]
GLEE-Net [20] 2024 TCSVT 61.9 54.0 75.8 45.9 55.7 77.6 92.9 60.0 65.5
SEV-Net [77] 2021 CVPR 55.3 53.1 61.5 53.6 38.2 71.6 95.7 41.5 58.8
MMA-Net [76] 2023 PRL 63.8 54.8 73.6 39.2 61.5 73.1 92.3 70.5 66.0
HMP-PS [78] 2021 CVPR 38.0 45.9 65.2 50.9 50.8 76.0 93.3 67.6 61.0
ISTR [80] 2021 ACM MM 47.5 53.3 64.4 51.8 44.4 74.7 92.1 60.7 61.1
MEGraph [56] 2022 IJCAI 52.5 45.7 76.1 51.8 46.5 76.1 92.9 57.6 62.4
AAR [79] 2023 TIP 62.4 53.6 71.5 39.0 48.8 76.1 91.3 70.6 64.2
SIMCLR [22] 2020 ICML 21.2 23.3 47.5 42.4 35.5 66.8 81.5 52.7 46.4
MoCo [32] 2020 CVPR 22.7 18.2 45.9 45.4 34.1 72.9 83.4 54.5 47.1
@@
BTFM [81] 2023 CVPR 41.5 44.9 60.3 51.5 50.3 70.4 91.3 55.3 58.2
SupHCL [49] 2022 ACM MM 52.5 58.8 70.0 53.5 51.4 73.1 95.6 58.0 64.1
KSRL [46] 2022 CVPR 60.4 59.2 67.5 52.7 51.5 76.1 91.3 57.7 64.5
CLEF [82] 2023 ICCV 64.3 61.8 68.4 49.0 55.2 72.9 89.9 57.0 64.8
AUNCE(Ours) - 61.8 58.9 74.9 49.7 56.2 73.5 92.1 64.2 66.4
(±$\pm 0.37$) (±$\pm 0.22$) (±$\pm 0.42$) (±$\pm 0.31$) (±$\pm 0.36$) (±$\pm 0.49$) (±$\pm 0.42$) (±$\pm 0.28$) (±$\pm 0.36$)
TABLE IV: Quantitative comparison results on GFT dataset, where the evaluation metric is F1-score(%).
Method Source AU Index Avg.
1 2 4 6 10 12 14 15 23 24
AlexNet [83] 2012 NIPS 44 46 2 73 72 82 5 19 43 42 42.8
LSVM [84] 2008 JMLR 38 32 13 67 64 78 15 29 49 44 42.9
TCAE [85] 2019 CVPR 43.9 49.5 6.3 71.0 76.2 79.5 10.7 28.5 34.5 41.7 44.2
CPAM [86] 2020 TBIOM 43.7 44.9 19.8 74.6 76.5 79.8 50.0 33.9 16.8 12.9 45.3
TAE [87] 2022 TPAMI 46.3 48.8 13.4 76.7 74.8 81.8 19.9 42.3 50.6 50.0 50.5
EAC-Net [7] 2018 TPAMI 15.5 56.6 0.1 81.0 76.1 84.0 0.1 38.5 57.8 51.2 46.1
ARL [15] 2019 TAC 51.9 45.9 13.7 79.2 75.5 82.8 0.1 44.9 59.2 47.5 50.1
@@
AAR [79] 2023 TIP 66.3 53.9 23.7 81.5 73.6 84.2 43.8 53.8 58.2 46.5 58.5
MAL [75] 2023 TAC 52.4 57.0 54.1 74.5 78.0 84.9 43.1 47.7 54.4 51.9 59.8
MoCo [32] 2020 CVPR 35.9 45.4 13.5 83.4 71.3 78.1 23.3 37.3 26.6 50.7 46.5
SimCLR [22] 2020 ICML 39.6 48.3 5.6 80.7 76.2 80.6 18.1 41.6 46.1 43.8 48.1
CLP [28] 2023 TIP 44.6 58.7 34.7 75.9 78.6 86.6 20.3 44.8 56.4 42.2 54.3
AUNCE(Ours) - 53.6 62.8 51.3 80.6 82.1 88.2 48.2 49.8 54.6 58.9 63.0
(±$\pm 0.33$) (±$\pm 0.18$) (±$\pm 0.28$) (±$\pm 0.24$) (±$\pm 0.41$) (±$\pm 0.11$) (±$\pm 0.35$) (±$\pm 0.23$) (±$\pm 0.42$) (±$\pm 0.35$) (±$\pm 0.29$)
TABLE V: Quantitative comparison results on Aff-Wild2 dataset, where the evaluation metric is F1-score(%).
Method Source AU Index Avg.
1 2 4 6 7 10 12 15 23 24 25 26
EAC-Net [7] 2018 TPAMI 49.6 33.7 55.6 66.4 82.3 81.4 76.9 11.8 12.5 12.2 93.7 26.8 50.2
ARL [15] 2019 TAC 59.2 48.2 54.9 70.0 83.4 80.3 72.0 0.1 0.1 17.3 93.0 37.5 51.3
@@
PASN [72] 2021 ICCV 65.7 64.2 66.5 76.6 74.7 72.7 78.6 18.5 10.6 55.1 80.7 41.7 58.8
AAR [79] 2023 TIP 65.4 57.9 59.9 73.2 84.6 83.2 79.9 21.8 27.4 19.9 94.5 41.7 59.1
AUNCE(Ours) - 68.8 67.6 64.7 74.8 88.6 78.4 81.3 26.9 36.7 18.6 96.1 43.8 62.2
(±$\pm 0.25$) (±$\pm 0.37$) (±$\pm 0.28$) (±$\pm 0.33$) (±$\pm 0.12$) (±$\pm 0.47$) (±$\pm 0.31$) (±$\pm 0.17$) (±$\pm 0.08$) (±$\pm 0.42$) (±$\pm 0.39$) (±$\pm 0.32$) (±$\pm 0.29$)
IV-B Comparison with State-of-the-Art Methods
IV-B1 Evaluation on BP4D
Since we use AU labels to assist the process of selecting positive sample pairs, we introduce five groups of supervised AU detection frameworks to compare with our method. The quantitative results of state-of-the-art methods from the past seven years for BP4D dataset are shown in Table II. AUNCE exhibits superior performance in AU detection when contrasted with alternative approaches.
Methods in the first group are based on feature extraction of regional area around AUs, including EAC-Net [7], ROI [4] and ARL [15]. Our approach achieves a notable improvement, with an increase in F1-score ranging from 5.0% to 10.2% over these methods. This advancement is primarily due to early feature extraction methods neglecting critical correlation information among AUs and failing to specifically address the class imbalance issue for each AU. Furthermore, the precision of the encoders utilized by these methods is generally lower.
@@ IV-B1 Evaluation on BP4D
Methods in the third group mainly utilize the AU semantic prior knowledge, including MMA-Net [76] and SEV-Net [77]. Our approach demonstrates an average improvement of 2.7% and 2.2%, respectively. This phenomenon can be attributed to the inherent limitations of these methods in effectively capturing and representing the distinct information encoded within Action Units (AUs) using simple semantic descriptions.
@@
Methods in the fourth group focus on capturing correlational information among AUs, including HMP-PS [78], AAR [79], ISTR [80] and MEGraph [56]. Our approach exhibits a performance advantage of approximately 0.6%-2.7% in F1-score. This underscores the efficacy of our supervised discriminative contrastive learning paradigm in effectively capturing differential information instead of pixel-level information. It is noteworthy that the encoder used in our method is only the same as the network MEGraph used for the first stage of training, with reduced parameters by nearly 7M. This highlights the applicability of AUNCE to more lightweight network structures and emphasizes that learning only the difference information among features is more potent than learning the pixel-level information of the entire image.
Methods in the last group are based on contrastive learning, including SIMCLR [22], MoCo [32], CLP [28], BTFM [81], SupHCL [49], KSRL [46] and CLEF [82]. When compared to these methods, our approach demonstrates a performance advantage of approximately 0.2% to 16.3% in F1-score. This improvement is attributed to the efficacy of our proposed negative sample re-weighting strategy and positive sample sampling strategy, specifically tailored for the AU detection task using static images.
IV-B2 Evaluation on DISFA
Comparable outcomes are evident in the DISFA dataset, as outlined in Table III, where AUNCE attains the highest overall performance in AU detection. Furthermore, a comparison between Table II and Table III reveals that recent methods such as SEV-Net, MEGraph, and BTFM demonstrate strong performance on the BP4D dataset but achieve only mediocre results on the DISFA dataset. This disparity suggests their limited generalizability. The more pronounced class imbalance issue in DISFA, as shown in Table I, likely contributes to this performance gap. In contrast, AUNCE consistently delivers robust performance on both BP4D and DISFA datasets and achieves more stable results across various AUs due to our proposed negative sample re-weighting strategy.
IV-B3 Evaluation on GFT
Table IV presents the F1-score results on the GFT dataset, where our AUNCE method demonstrates significantly superior performance compared to previous approaches. Unlike BP4D and DISFA, which primarily consist of near-frontal face images, GFT includes many images captured under out-of-plane poses. In this challenging scenario, other methods such as EAC-Net, ARL and CLP so on struggle to perform well. In contrast, AUNCE achieves a commendable average F1-score of 63.0%.
IV-B4 Evaluation on Aff-Wild2
[…]
IV-C Ablation Study
IV-C1 Compared with weighted cross-entropy loss
@@
TABLE VI: Comparison of the weighted cross-entropy (WCE) loss and AUNCE with different encoders on BP4D dataset.
Backbone Source Loss F1-Score(%) Accuracy(%)
ResNet18 [88] 2016CVPR WCE 52.3 73.8
AUNCE $\textbf{54.9}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2.6\uparrow}}$ $\textbf{74.3}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.5\uparrow}}$
ResNet50 [88] 2016CVPR WCE 56.2 75.5
AUNCE $\textbf{59.7}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3.5\uparrow}}$ $\textbf{77.0}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.5\uparrow}}$
ROI [4] 2017CVPR WCE 56.4 75.9
AUNCE $\textbf{58.8}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2.4\uparrow}}$ $\textbf{76.7}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.8\uparrow}}$
ARL [15] 2019TAC WCE 61.1 78.2
AUNCE $\textbf{62.9}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.8\uparrow}}$ $\textbf{79.1}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.9\uparrow}}$
[…]
AUNCE $\textbf{64.1}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.7\uparrow}}$ $\textbf{80.2}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.6\uparrow}}$
MEGraph [56] 2022IJCAI WCE 63.6 80.0
(Onestage) AUNCE $\textbf{66.1}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2.5\uparrow}}$ $\textbf{80.4}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.4\uparrow}}$
Due to the primary innovation of this paper being the introduction of a novel contrastive loss, AUNCE, which is compatible with various AU detection encoders, we conducted a comparative analysis between the proposed AUNCE loss and the traditional weighted cross-entropy loss (WCE) across multiple AU detection encoders, as detailed in Table VI. The findings reveal that AUNCE demonstrates superior performance in AU detection. Notably, the MEGraph encoder achieved the best results, with the F1-score on the BP4D dataset improving from 63.6% to 66.1%. This underscores the effectiveness of capturing unique features of different samples over focusing solely on pixel-level information. Furthermore, we optimized the two-stage training process of MEGraph into a single-stage process, significantly reducing both the training time and the number of parameters required for the model.
TABLE VII: Quantitative results of ablation study on BP4D dataset.
Model BL wi PS NS AU Index F1-Score(%) Accuracy(%)
1 2 4 6 7 10 12 14 15 17 23 24
A $\checkmark$ 39.8 44.3 58.8 76.0 77.7 76.9 88.0 64.5 49.0 59.0 43.5 43.4 60.1 79.1
B $\checkmark$ $\checkmark$ 40.1 43.0 58.9 75.4 79.3 78.8 87.8 64.1 48.7 61.0 46.9 47.5 61.00.9↑ 79.30.2↑
C $\checkmark$ $\checkmark$ $\checkmark$ 51.1 45.6 58.3 77.5 77.0 84.5 89.0 63.7 54.6 65.5 46.8 53.3 63.83.7↑ 79.80.7↑
D $\checkmark$ $\checkmark$ $\checkmark$ 50.2 43.0 58.7 77.9 77.3 83.7 88.7 64.2 53.9 64.9 50.0 56.4 64.14.0↑ 79.90.8↑
E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 53.6 49.8 61.6 78.4 78.8 84.7 89.6 67.4 55.1 65.4 50.9 58.0 $\textbf{66.1}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}6.0\uparrow}}$ $\textbf{80.4}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.3\uparrow}}$
IV-C2 Evaluating other components
We further evaluate the effectiveness of weight coeffcient $\omega_{i}$, negative sample re-weighting strategy and positive sample samplying strategy. Quantitative results of our method with the above component combinations are summarized in Table VII. The baseline (BL) is rooted in SIMCLR [22], but positive sample pairs are changed to images with the highest similarity to original images. Other components are the weight coefficient for both WCE and AUNCE ($w_{i}$), the positive sample sampling strategy (PS), and the negative sample re-weighting strategy (NS).
Weight coefficient $w_{i}$: The weights $w_{i}$ have been pointed out as a fixed value in Section. III-A, which achieve an increase of 0.9% in F1-score and 0.2% in Accuracy. For an AU, its activation indicator is a binary variable, and the corresponding information content can be measured using entropy: $-\log(p_{i})$. Therefore, AUs with lower activation probabilities carry more information and should be assigned higher weights. This forms the theoretical basis for our weighting strategy.
Positive sample sampling strategy: As outlined in Table VII, the positive sample sampling strategy contributes to a 2.8% increase in F1-score and a 0.5% improvement in Accuracy compared to model B. This validates that selecting different types of positive sample pairs with appropriate probabilities can mitigate issues arising from AU labels tainted by noise and errors. The strategy achieves this by incorporating three types of positive sample pairs: the self-supervised signal pair, the highest similarity pair, and the mixed positive features pair. These pairs provide the models with diverse information from various views, thereby enhancing the generalization ability of the contrastive learning framework.
Negative sample re-weighting strategy: The negative sample re-weighting strategy to minority AUs results in a 3.1% increase in F1-score and a 0.6% increase in Accuracy compared to model B. This illustrates that the strategy can change the backpropagation gradient of minority and majority class samples, facilitating faster label category gradient updates with a smaller number of samples. Additionally, model E achieves an average increase of 2.3% in F1-score compared to model C, further supporting the efficacy of the re-weighting strategy.
Figure 4: T-SNE visualization of feature representations on BP4D and DISFA datasets. Colors indicates whether AU1 exists. Top row: (a) Representations optimized by WCE on BP4D dataset, and (b)-(f) Visualizations of ablation study on BP4D dataset, respectively represent model A-E. Bottom row: (g) Representations optimized by AUNCE on DISFA dataset, and (h)-(l) Visualizations of ablation study on DISFA dataset, respectively represent model A-E.
We further show the visualization using T-SNE dimensionality reduction of the learned feature representations optimized by AUNCE and WCE to qualitatively assess the efficacy of our proposed discriminative contrastive learning paradigm, as illustrated in Fig. 4. For each dataset, a random sample of 200 images is utilized for illustration purposes. A comparison between Fig. 4(a), Fig. 4(f) against Fig. 4(g), Fig. 4(l) reveals that the learned representations optimized by AUNCE exhibit greater separability than those optimized by WCE. Additionally, Fig. 4(b)-(f) and Fig. 4(g)-(l) qualitatively showcase the effectiveness of each component in our ablation study.